OpenStudy (anonymous):

Calculate the equilibrium constant for the oxidation of I-(aq) to IO3-(aq) by O2 at 25⁰C in an aqueous acid solution, according to the following chemical equation 2 I-(aq) + 3 O2(g) D 2 IO3-(aq) Data: E0 (V) 2 IO3-(aq) + 12 H+(aq) + 12 e- → 2 I-(aq) + 6 H2O(l) 1.09 O2(g) + 4 H+(aq) + 4 e- → 2 H2O(l) 1.23

OpenStudy (aaronq):

(1) 2 IO3-(aq) + 12 H+(aq) + 12 e- → 2 I-(aq) + 6 H2O(l) 1.09 (2) O2(g) + 4 H+(aq) + 4 e- → 2 H2O(l) 1.23 ----------------------------------------------------------- (2) multiply by 3 (2) 3O2(g) + 12 H+(aq) + 12 e- → 6 H2O(l) 1.23 ----------------------------------------------------------- Reverse (1) (1) 2 I-(aq) + 6 H2O(l) → 2 IO3-(aq) + 12 H+(aq) + 12 e- 1.09 ------------------------------------------------------------- (1) 2 I-(aq) + 6 H2O(l) → 2 IO3-(aq) + 12 H+(aq) + 12 e- 1.09 (2) 3O2(g) + 12 H+(aq) + 12 e- → 6 H2O(l) 1.23 cancel stuff out: 2 I-(aq) + 3O2(g) →2 IO3-(aq) E = cathode-anode = 1.23-1.09 = 0.24 V then theres 2 ways to find K: (1) delta G= -nFE = -(12)(96,485.33)(0.24) = - 277877.75 delta G= -RTlnK = - 277877.75 = -(8.314)(298.15)lnK K = 4.84*10^48 (2) Nernst equation: at 25 degrees, its simplified to: Ecell=E(knot)-(0.0592/n)(logK) at equilibrium Ecell=0, so E(knot)=(0.0592/n)(logK) 0.24 = (0.0592/12)logK K= 4.45*10^48 there is a slight difference in the K values because of the simplification of the nernst equation.

OpenStudy (anonymous):

intense, thank you!

OpenStudy (anonymous):

What is delta G by the way?

OpenStudy (aaronq):

haha i know, electrochemistry looks intense, but it isn't really. \[\Delta G \] is gibbs free energy

OpenStudy (anonymous):

oooo we never talked about gibbs free energy... :o

OpenStudy (anonymous):

is it used in any other places other than for those two equation you used, -nFE and -RTlnK?

OpenStudy (aaronq):

yep, it's used in physics, and thermodynamics

OpenStudy (aaronq):

a big one is \[\Delta G = \Delta H - T \Delta S\]

OpenStudy (anonymous):

o.O Could you tell me this Gibbs Free energy really represents... ?

OpenStudy (anonymous):

some energy thatès not directly equal to entropy or enthalpy?

OpenStudy (aaronq):

it's defined as the free energy in the system, it can be seen as the stability of a molecule but its much more than that. I mean this is some deep stuff that I'm not comfortable with enough to explain it, but if you watch this vid, you might understand it more http://www.youtube.com/watch?v=DPjMPeU5OeM

OpenStudy (anonymous):

thank you!

OpenStudy (aaronq):

no prob dude